# Beyond Random: Automatic Inner-loop Optimization in Dataset Distillation

### v1

@@ 1 Introduction
The unprecedented success of deep learning DBLP:conf/icml/IoffeS15 ; DBLP:journals/pami/LeiT24 has been fundamentally driven by the availability of large-scale datasets and computational resources. However, this data-centric paradigm poses scalability challenges, particularly in storage efficiency, computational overhead, and environmental sustainability DBLP:conf/nips/CuiWSH22 ; DBLP:journals/corr/SimonyanZ14a ; DBLP:conf/mm/LiZ0XLQ24 . These limitations have catalyzed the emergence of dataset distillation DBLP:journals/corr/abs-1811-10959 ; DBLP:conf/iclr/ZhaoMB21 ; DBLP:conf/ijcai/GengCWWSMZR23 ; DBLP:journals/pami/YuLW24 ; DBLP:conf/aaai/LiZDXQ25 (DD) as a transformative approach to data-efficient learning. By synthesizing compact surrogate datasets that preserve the essential characteristics of their original counterparts, DD enables efficient model training without compromising performance.
Dataset distillation is inherently formulated as a bilevel optimization problem DBLP:conf/cvpr/SunSYL24 ; feng2024embarrassingly , consisting of inner-loop and outer-loop optimization. The outer-loop optimizes the synthetic dataset to minimize discrepancies between models trained on distilled and original data, and the inner-loop simulates the training dynamics of a neural network on the distilled dataset feng2024embarrassingly . Previous DD methods focus on approximating the outer-loop to directly align the final performance metric DBLP:conf/iclr/ZhaoMB21 ; DBLP:conf/icml/ZhaoB21 ; DBLP:conf/icml/LooHLR23 ; DBLP:conf/cvpr/Cazenavette00EZ22b ; DBLP:conf/icml/CuiWSH23 ; DBLP:conf/icml/LeeC24 ; DBLP:conf/iclr/Guo0CLZ024 ; DBLP:conf/eccv/LiuGCTS24 . However, outer-loop methods typically rely on surrogate objectives or heuristic measures to indirectly capture the effects of inner-loop training, thereby failing to accurately reflect the model’s training dynamics. Popular inner-loop optimization techniques, such as Backpropagation Through Time (BPTT) DBLP:journals/pieee/Werbos90 and its truncated variants (T-BPTT DBLP:conf/aistats/ShabanCHB19 , RaT-BPTT feng2024embarrassingly ), adopt rigid truncation strategies to partially address performance bottlenecks. Nevertheless, these methods uniformly apply fixed or random truncation across all training stages, disregarding the distinct learning patterns of neural networks: early stages prioritize simple patterns and later stages refine complex features DBLP:conf/icml/ArpitJBKBKMFCBL17 . Consequently, the rigid truncation of learning trajectories limits model performance.
@@
To address these limitations, we posit that automatic adjustment of truncation strategies across distinct training stages could better align with the intrinsic learning dynamics of deep neural networks (DNNs) DBLP:conf/icml/ArpitJBKBKMFCBL17 . We verify this perspective through a controlled variable experiment, which partitions the training process into three stages, and selects preliminary or post timesteps unfolding in each epoch as the truncation position. Our experimental analysis reveals three critical findings: (1) early-stage truncation of preliminary timesteps yields a 2.9% improvement in test accuracy (Fig. 1a), (2) middle-stage performance demonstrates negligible sensitivity to truncation position ($\Delta<0.3$%, Fig. 1b), and (3) late-stage prioritization of post timesteps enhances accuracy by 1.8% (Fig. 1c). These results substantiate the suboptimality of random truncation approaches, and motivate the formulation of a secondary perspective: adaptive window sizing could simultaneously optimize computational efficiency and gradient preservation. Subsequent experiments with fixed optimal truncation positions (Fig. 1d-f) demonstrate that a larger window in the early stage achieves a 2.5% accuracy gain, while the middle- and late-stage variations show insignificant impacts ($\Delta<0.2$%). These hypotheses and their validation are detailed in Section 4.1.
Based on the above observation, we propose an effective DD method called Automatic Truncated Backpropagation Through Time (AT-BPTT). This method integrates three key mechanisms: dynamic truncation position, adaptive window size and low-rank Hessian approximation. Dynamic truncation position employs gradient magnitudes to probabilistically determine better truncation timesteps across different training stages. Adaptive window size leverages the magnitudes of gradient variation to modulate the truncation window, ensuring that critical gradient information is retained when fluctuations are high. Low-rank Hessian approximation addresses the computational time and storage cost issues of Hessian matrix calculation in second-order optimization. Extensive experiments illustrate that AT-BPTT achieves state-of-the-art (SOTA) performance on CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny and ImageNet-1K DBLP:journals/ijcv/RussakovskyDSKS15 , outperforming the leading inner-loop method by an average of 6.16% while delivering a 3.9 × faster training speed and a 63% memory reduction.
@@ 2 Related Work
Dataset distillation is proposed with the aim of generating a compact synthetic dataset that effectively substitutes for the original dataset. Dataset distillation can be formulated as a bilevel optimization problem, with current mainstream approaches broadly categorized into outer-loop optimization methods (data matching methods) and inner-loop optimization methods (meta-learning methods). Outer-loop optimization methods achieve dataset distillation by aligning surrogate objectives between the original and synthetic datasets. Gradient matching DBLP:conf/icml/ZhaoB21 aligns the gradients of synthetic and original data during training to mimic the optimization behavior of the original dataset. Trajectory matching DBLP:conf/cvpr/Cazenavette00EZ22b ; DBLP:conf/icml/CuiWSH23 ; DBLP:conf/iclr/Guo0CLZ024 ; DBLP:conf/eccv/LiuGCTS24 ; DBLP:journals/corr/abs-2503-18872 aligns parameter trajectories throughout the training process, overcoming the short-term matching limitations of gradient matching. Distribution matching DBLP:conf/cvpr/WangZPZYWHBWY22 ; DBLP:conf/wacv/ZhaoB23 ; DBLP:journals/corr/abs-2502-20653 aligns distributions in the feature space, circumventing the complexity of bilevel optimization. However, outer-loop methods rely on surrogate models to indirectly capture the effects of inner-loop training, failing to accurately reflect the true training dynamics of the model.
Inner-loop optimization methods treat the synthetic dataset as hyperparameters, minimizing the risk of models trained on synthetic data on the target dataset. A prominent inner-loop approach is Backpropagation Through Time DBLP:journals/pieee/Werbos90 , which simulates the model training process by performing multi-step gradient descent in the inner loop and optimizes the synthetic data in the outer loop to achieve dataset distillation. Subsequently, T-BPTT DBLP:conf/aistats/ShabanCHB19 improves distillation performance by truncating the unrolled time steps. Recently, RaT-BPTT feng2024embarrassingly employs a random truncation strategy to enhance BPTT’s distillation performance, partially reducing computational costs. However, its random truncation strategy fails to align with the learning characteristics of deep neural networks, limiting its performance.
Besides optimization-based techniques, there are several other DD methods. Diffusion-based methods DBLP:conf/cvpr/SuHG0T24 ; DBLP:conf/iclr/ChenDH0Z025 generate compact samples that capture key features of the original dataset by optimizing representations in the latent space or directly adjusting the diffusion process, excelling in high-quality image generation and high-resolution tasks. Decoupled optimization methods DBLP:conf/nips/YinXS23 ; DBLP:journals/corr/abs-2411-19946 ; DBLP:journals/corr/abs-2503-18267 decompose the complex distillation process into independently optimizable subtasks, reducing computational complexity and enhancing the capability to handle large-scale datasets.
@@ 3 Preliminary
Dataset distillation feng2024embarrassingly aims to compress a large training set $\mathcal{D}$ into a smaller set $\mathcal{S}$ such that training on $\mathcal{S}$ achieves comparable performance to training on $\mathcal{D}$. The process involves a bilevel formulation:
@@
Backpropagation Through Time (BPTT). BPTT DBLP:journals/pieee/Werbos90 is the standard method for solving bilevel optimization in reverse mode. When $\mathcal{A}$ follows gradient descent with learning rate $\alpha$, the meta-gradient with respect to every unrolling learning trajectory is computed using the chain rule:
@@
Truncated BPTT (T-BPTT). To alleviate the memory burden, T-BPTT DBLP:conf/aistats/ShabanCHB19 propagates gradients through a smaller unrolling window of $M$ steps instead of the full trajectory:
@@
Random Truncated BPTT (RaT-BPTT). RaT-BPTT feng2024embarrassingly extends T-BPTT by randomly positioning the truncated window along every unrolling learning trajectory in the training process. The meta-gradient of RaT-BPTT with random truncation at position $N$ is:
@@ 4.1 Hypothesis Verification
We challenge the practice of applying random truncation uniformly across the entire training process in RaT-BPTT feng2024embarrassingly based on two key observations. First, it is well established that deep neural networks (DNNs) tend to learn simple patterns during the early training stages before progressively acquiring more complex patterns DBLP:conf/icml/ArpitJBKBKMFCBL17 . Second, T-BPTT DBLP:conf/aistats/ShabanCHB19 exclusively utilizes the last TM+1T-M+1 timesteps of each epoch, while the impact of restricting updates to early timesteps remains unexplored. These observations motivate our Hypothesis I: Might sampling specific timesteps from distinct stages lead to better performance, rather than uniformly applying random truncation?
To verify this hypothesis, we partition the epochs throughout the training process into three equal stages (early, middle, and late stages) while bisecting the timesteps within the unfolding learning trajectory for every epoch (preliminary and post phases). Our validation involves controlled experiments where we randomly truncate either preliminary or post timesteps during the early stage, maintaining standard RaT-BPTT implementation in subsequent two stages. Similar validations are implemented on the middle and late training stages. The experiments are conducted on the CIFAR-10 krizhevsky2009learning dataset using the ConvNets architecture. Our experimental results shown in Fig. 1a-c reveal three key observations: (1) preliminary phase truncation in early stage enhances validation accuracy by an average of 2.9%; (2) middle stage shows negligible performance variance ($\Delta<0.5$%) between phase choices; (3) post phase selection in late stage demonstrates 1.8% accuracy improvement. These findings not only validate our initial hypothesis but also raise a new Hypothesis II: Can dynamically adjusting the truncation window size further enhance distillation performance?
@@ 4.2 Automatic Truncated BPTT
Dynamic Truncation Position. The experimental validation of Hypothesis I raises a question: How to establish an evaluation metric that enables staged truncation of timesteps during different training stages? Our analysis begins with the gradient accumulation mechanism expressed in Eqn. 4, which reveals that parameter updates represent cumulative contributions from all timestep gradients, with each timestep’s influence modulated by its temporal position and subsequent gradient dynamics. To quantify these temporal effects, we record gradient magnitudes across all timesteps during training iterations, with averages visualized in Fig. 2. It is observed that gradient updates are initially large but gradually diminish as timesteps progress. Interestingly, this observation and the verification of Hypothesis I align with Arpit et al. DBLP:conf/icml/ArpitJBKBKMFCBL17 ’s seminal work on deep learning dynamics: DNNs tend to learn simple, easily identifiable patterns during the early stages of training before progressively shifting toward more complex and fine-grained patterns in later stages. This is attributable to that simple patterns often dominate the data and exhibit high saliency, allowing large gradient updates in the preliminary phase to align effectively with these patterns, leading to rapid performance improvement in the early stage. Although the gradients of timesteps in the post phase are smaller, they play a crucial role in fine-tuning the network, ultimately leading to significant performance gains in the late stage.
@@ 5.1 Experimental Setup
Dataset. We adhere to the conventional procedure adopted in dataset distillation feng2024embarrassingly . We select three standard datasets: CIFAR-10 krizhevsky2009learning (10 classes, $\times32$), CIFAR-100 krizhevsky2009learning (100 classes, $\times32$), and Tiny-ImageNet le2015tiny (200 classes, $\times64$). To further show the effectiveness of AT-BPTT on high-resolution images, we scale up the dataset to ImageNet-1K DBLP:journals/ijcv/RussakovskyDSKS15 (1,000 classes, $\times224$). For CIFAR-10 and CIFAR-100, we distill datasets with 1, 10, and 50 images per class (IPC = 1, 10, 50), while for Tiny-ImageNet and ImageNet-1K, we use 1 and 10 images per class (IPC = 1, 10).
@@
Implementation Details. Following Rat-BPTT feng2024embarrassingly , we employ standardized convolutional neural network DBLP:conf/cvpr/GidarisK18 (CNN) architectures with depth adapted to dataset specifications: a 3-layer CNN (Conv-3) for CIFAR-10/CIFAR-100 synthetic datasets, and a 4-layer CNN (Conv-4) for Tiny-ImageNet and high-resolution ImageNet-1K datasets. More hyperparameter settings are detailed in Appendix B.
@@ 5.2 Comparison with Previous Methods
Tab. 1 demonstrates that the proposed AT-BPTT framework establishes new state-of-the-art performance on three benchmark datasets (CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , and Tiny-ImageNet le2015tiny ), outperforming existing inner-loop and outer-loop optimization approaches by a significant margin. Most remarkably, AT-BPTT achieves an average accuracy gain of 6.16% over RaT-BPTT, the SOTA inner-loop DD method. This substantial improvement in low-resolution image distillation stems from the combination of our novel dynamic truncation strategy and adaptive window size mechanism, which collectively enhance the model’s capacity to capture multi-level feature correlations through dynamically adjusted learning trajectories. The visualizations of these synthetic datasets are shown in Appendix F.
@@
Table 1: Comparison with the SOTA baseline dataset distillation methods. Following previous methods, the ConvNet used for distillation are Conv-3 on CIFAR-10 krizhevsky2009learning and CIFAR-100 krizhevsky2009learning , Conv-4 on Tiny-ImageNet le2015tiny and ImageNet-1K DBLP:journals/ijcv/RussakovskyDSKS15 . Each reported result is the average of 5 experiments. Entries with “-” are absent due to scalability problems.
@@
DSA DBLP:conf/iclr/ZhaoMB21 28.8±0.7 52.1±0.5 60.6±0.5 13.9±0.3 32.3±0.3 42.8±0.4 6.6±0.2 14.4±2.0 1.1±0.7 3.2±0.3
CAFE DBLP:conf/cvpr/WangZPZYWHBWY22 30.3±1.1 46.3±0.6 55.5±0.6 12.9±0.3 27.8±0.3 37.9±0.3 - - - -
MTT DBLP:conf/cvpr/Cazenavette00EZ22b 46.2±0.8 65.4±0.7 71.6±0.2 24.3±0.3 39.7±0.4 47.7±0.2 8.8±0.3 23.2±0.2 - -
TESLA DBLP:conf/icml/CuiWSH23 48.5±0.8 66.4±0.8 72.6±0.7 24.8±0.4 41.7±0.3 47.9±0.3 - - 7.7±1.0 17.8±0.5
DATM DBLP:conf/iclr/Guo0CLZ024 46.9±0.5 66.8±0.2 76.1±0.3 27.9±0.2 47.2±0.4 55.0±0.2 17.1±0.3 31.1±0.3 - -
ATT DBLP:conf/eccv/LiuGCTS24 48.3±1.0 67.7±0.6 74.5±0.9 26.1±0.5 44.2±0.8 51.2±0.2 11.0±0.4 25.8±0.7 4.7±1.4 8.7±1.0
MCT DBLP:journals/corr/abs-2406-19827 48.5±0.2 66.0±0.3 72.3±0.3 24.5±0.5 42.5±0.5 46.8±0.2 9.6±0.5 22.6±0.8 - -
NCFM DBLP:journals/corr/abs-2502-20653 49.5±0.3 71.8±0.3 77.4±0.3 34.4±0.5 48.7±0.3 54.7±0.2 18.2±0.5 26.8±0.6 - -
@@
BPTT DBLP:journals/pieee/Werbos90 49.1±0.6 62.4±0.4 70.5±0.4 21.3±0.6 34.7±0.5 - - - 1.1±0.7 2.3±0.9
FRePO DBLP:conf/nips/ZhouNB22 45.6±0.1 63.5±0.1 70.7±0.1 26.3±0.1 41.3±0.1 41.5±0.1 16.9±0.1 22.4±0.1 - -
RCIG DBLP:conf/icml/LooHLR23 49.6±1.2 66.8±0.3 - 35.5±0.7 - - 22.4±0.3 - - -
RaT-BPTT feng2024embarrassingly 53.2±0.7 69.4±0.4 75.3±0.3 35.3±0.4 47.5±0.2 50.6±0.2 20.1±0.3 24.4±0.2 5.2±1.1 13.0±0.9
Teddy DBLP:conf/eccv/YuLYW24 30.1±1.4 53.0±0.5 66.1±0.4 13.5±0.4 33.4±0.7 49.4±0.5 - - - 34.1±0.8

### v2

@@ 1 Introduction
The unprecedented success of deep learning Ioffe and Szegedy (2015); Lei and Tao (2024) has been fundamentally driven by the availability of large-scale datasets and computational resources. However, this data-centric paradigm poses scalability challenges, particularly in storage efficiency, computational overhead, and environmental sustainability Cui et al. (2022); Simonyan and Zisserman (2015); Li et al. (2024a). These limitations have catalyzed the emergence of dataset distillation Wang et al. (2018); Zhao et al. (2021); Geng et al. (2023); Yu et al. (2024a); Li et al. (2025) (DD) as a transformative approach to data-efficient learning. By synthesizing compact surrogate datasets that preserve the essential characteristics of their original counterparts, DD enables efficient model training without compromising performance.
Dataset distillation is inherently formulated as a bilevel optimization problem Sun et al. (2024); Feng et al. (2024), consisting of inner-loop and outer-loop optimization. The outer-loop optimizes the synthetic dataset to minimize discrepancies between models trained on distilled and original data, and the inner-loop simulates the training dynamics of a neural network on the distilled dataset Feng et al. (2024). Previous DD methods focus on approximating the outer-loop to directly align the final performance metric Zhao et al. (2021); Zhao and Bilen (2021); Loo et al. (2023); Cazenavette et al. (2022); Cui et al. (2023); Lee and Chung (2024); Guo et al. (2024); Liu et al. (2024). However, outer-loop methods typically rely on surrogate objectives or heuristic measures to indirectly capture the effects of inner-loop training, thereby failing to accurately reflect the model’s training dynamics. Popular inner-loop optimization techniques, such as Backpropagation Through Time (BPTT) Werbos (1990) and its truncated variants (T-BPTT Shaban et al. (2019), RaT-BPTT Feng et al. (2024)), adopt rigid truncation strategies to partially address performance bottlenecks. Nevertheless, these methods uniformly apply fixed or random truncation across all training stages, disregarding the distinct learning patterns of neural networks: early stages prioritize simple patterns and later stages refine complex features Arpit et al. (2017). Consequently, the rigid truncation of learning trajectories limits model performance.
@@
To address these limitations, we posit that automatic adjustment of truncation strategies across distinct training stages could better align with the intrinsic learning dynamics of deep neural networks (DNNs) Arpit et al. (2017). We verify this perspective through a controlled variable experiment, which partitions the training process into three stages, and selects preliminary or post timesteps unfolding in each epoch as the truncation position. Our experimental analysis reveals three critical findings: (1) early-stage truncation of preliminary timesteps yields a 2.9% improvement in test accuracy (Fig. 1a), (2) middle-stage performance demonstrates negligible sensitivity to truncation position ($\Delta<0.3$%, Fig. 1b), and (3) late-stage prioritization of post timesteps enhances accuracy by 1.8% (Fig. 1c). These results substantiate the suboptimality of random truncation approaches, and motivate the formulation of a secondary perspective: adaptive window sizing could simultaneously optimize computational efficiency and gradient preservation. Subsequent experiments with fixed optimal truncation positions (Fig. 1d-f) demonstrate that a larger window in the early stage achieves a 2.5% accuracy gain, while the middle- and late-stage variations show insignificant impacts ($\Delta<0.2$%). These hypotheses and their validation are detailed in Section 4.1.
Based on the above observation, we propose an effective DD method called Automatic Truncated Backpropagation Through Time (AT-BPTT). This method integrates three key mechanisms: dynamic truncation position, adaptive window size and low-rank Hessian approximation. Dynamic truncation position employs gradient magnitudes to probabilistically determine better truncation timesteps across different training stages. Adaptive window size leverages the magnitudes of gradient variation to modulate the truncation window, ensuring that critical gradient information is retained when fluctuations are high. Low-rank Hessian approximation addresses the computational time and storage cost issues of Hessian matrix calculation in second-order optimization. Extensive experiments illustrate that AT-BPTT achieves state-of-the-art (SOTA) performance on CIFAR-10 Krizhevsky et al. (2009), CIFAR-100 Krizhevsky et al. (2009), Tiny-ImageNet Le and Yang (2015) and ImageNet-1K Russakovsky et al. (2015), outperforming the leading inner-loop method by an average of 6.16% while delivering a 3.9 × faster training speed and a 63% memory reduction.
@@ 2 Related Work
Dataset distillation is proposed with the aim of generating a compact synthetic dataset that effectively substitutes for the original dataset. Dataset distillation can be formulated as a bilevel optimization problem, with current mainstream approaches broadly categorized into outer-loop optimization methods (data matching methods) and inner-loop optimization methods (meta-learning methods). Outer-loop optimization methods achieve dataset distillation by aligning surrogate objectives between the original and synthetic datasets. Gradient matching Zhao and Bilen (2021) aligns the gradients of synthetic and original data during training to mimic the optimization behavior of the original dataset. Trajectory matching Cazenavette et al. (2022); Cui et al. (2023); Guo et al. (2024); Liu et al. (2024); Chen et al. (2025b) aligns parameter trajectories throughout the training process, overcoming the short-term matching limitations of gradient matching. Distribution matching Wang et al. (2022); Zhao and Bilen (2023); Wang et al. (2025c) aligns distributions in the feature space, circumventing the complexity of bilevel optimization. However, outer-loop methods rely on surrogate models to indirectly capture the effects of inner-loop training, failing to accurately reflect the true training dynamics of the model.
Inner-loop optimization methods treat the synthetic dataset as hyperparameters, minimizing the risk of models trained on synthetic data on the target dataset. A prominent inner-loop approach is Backpropagation Through Time Werbos (1990), which simulates the model training process by performing multi-step gradient descent in the inner loop and optimizes the synthetic data in the outer loop to achieve dataset distillation. Subsequently, T-BPTT Shaban et al. (2019) improves distillation performance by truncating the unrolled time steps. Recently, RaT-BPTT Feng et al. (2024) employs a random truncation strategy to enhance BPTT’s distillation performance, partially reducing computational costs. However, its random truncation strategy fails to align with the learning characteristics of deep neural networks, limiting its performance.
Besides optimization-based techniques, there are several other DD methods. Diffusion-based methods Su et al. (2024b); Chen et al. (2025a) generate compact samples that capture key features of the original dataset by optimizing representations in the latent space or directly adjusting the diffusion process, excelling in high-quality image generation and high-resolution tasks. Decoupled optimization methods Yin et al. (2023); Shen et al. (2024); Tran et al. (2025) decompose the complex distillation process into independently optimizable subtasks, reducing computational complexity and enhancing the capability to handle large-scale datasets.
@@ 3 Preliminary
Dataset distillation Feng et al. (2024) aims to compress a large training set $\mathcal{D}$ into a smaller set $\mathcal{S}$ such that training on $\mathcal{S}$ achieves comparable performance to training on $\mathcal{D}$. The process involves a bilevel formulation:
@@
Backpropagation Through Time (BPTT). BPTT Werbos (1990) is the standard method for solving bilevel optimization in reverse mode. When $\mathcal{A}$ follows gradient descent with learning rate $\alpha$, the meta-gradient with respect to every unrolling learning trajectory is computed using the chain rule:
@@
Truncated BPTT (T-BPTT). To alleviate the memory burden, T-BPTT Shaban et al. (2019) propagates gradients through a smaller unrolling window of $M$ steps instead of the full trajectory:
@@
Random Truncated BPTT (RaT-BPTT). RaT-BPTT Feng et al. (2024) extends T-BPTT by randomly positioning the truncated window along every unrolling learning trajectory in the training process. The meta-gradient of RaT-BPTT with random truncation at position $N$ is:
@@ 4.1 Hypothesis Verification
We challenge the practice of applying random truncation uniformly across the entire training process in RaT-BPTT Feng et al. (2024) based on two key observations. First, it is well established that deep neural networks (DNNs) tend to learn simple patterns during the early training stages before progressively acquiring more complex patterns Arpit et al. (2017). Second, T-BPTT Shaban et al. (2019) exclusively utilizes the last TM+1T-M+1 timesteps of each epoch, while the impact of restricting updates to early timesteps remains unexplored. These observations motivate our Hypothesis I: Might sampling specific timesteps from distinct stages lead to better performance, rather than uniformly applying random truncation?
To verify this hypothesis, we partition the epochs throughout the training process into three equal stages (early, middle, and late stages) while bisecting the timesteps within the unfolding learning trajectory for every epoch (preliminary and post phases). Our validation involves controlled experiments where we randomly truncate either preliminary or post timesteps during the early stage, maintaining standard RaT-BPTT implementation in subsequent two stages. Similar validations are implemented on the middle and late training stages. The experiments are conducted on the CIFAR-10 Krizhevsky et al. (2009) dataset using the ConvNets architecture. Our experimental results shown in Fig. 1a-c reveal three key observations: (1) preliminary phase truncation in early stage enhances validation accuracy by an average of 2.9%; (2) middle stage shows negligible performance variance ($\Delta<0.5$%) between phase choices; (3) post phase selection in late stage demonstrates 1.8% accuracy improvement. These findings not only validate our initial hypothesis but also raise a new Hypothesis II: Can dynamically adjusting the truncation window size further enhance distillation performance?
@@ 4.2 Automatic Truncated BPTT
Dynamic Truncation Position. The experimental validation of Hypothesis I raises a question: How to establish an evaluation metric that enables staged truncation of timesteps during different training stages? Our analysis begins with the gradient accumulation mechanism expressed in Eqn. 4, which reveals that parameter updates represent cumulative contributions from all timestep gradients, with each timestep’s influence modulated by its temporal position and subsequent gradient dynamics. To quantify these temporal effects, we record gradient magnitudes across all timesteps during training iterations, with averages visualized in Fig. 2. It is observed that gradient updates are initially large but gradually diminish as timesteps progress. Interestingly, this observation and the verification of Hypothesis I align with Arpit et al. Arpit et al. (2017)’s seminal work on deep learning dynamics: DNNs tend to learn simple, easily identifiable patterns during the early stages of training before progressively shifting toward more complex and fine-grained patterns in later stages. This is attributable to that simple patterns often dominate the data and exhibit high saliency, allowing large gradient updates in the preliminary phase to align effectively with these patterns, leading to rapid performance improvement in the early stage. Although the gradients of timesteps in the post phase are smaller, they play a crucial role in fine-tuning the network, ultimately leading to significant performance gains in the late stage.
@@ 5.1 Experimental Setup
Dataset. We adhere to the conventional procedure adopted in dataset distillation Feng et al. (2024). We select three standard datasets: CIFAR-10 Krizhevsky et al. (2009) (10 classes, $\times32$), CIFAR-100 Krizhevsky et al. (2009) (100 classes, $\times32$), and Tiny-ImageNet Le and Yang (2015) (200 classes, $\times64$). To further show the effectiveness of AT-BPTT on high-resolution images, we scale up the dataset to ImageNet-1K Russakovsky et al. (2015) (1,000 classes, $\times224$). For CIFAR-10 and CIFAR-100, we distill datasets with 1, 10, and 50 images per class (IPC = 1, 10, 50), while for Tiny-ImageNet and ImageNet-1K, we use 1 and 10 images per class (IPC = 1, 10).
@@
Implementation Details. Following Rat-BPTT Feng et al. (2024), we employ standardized convolutional neural network Gidaris and Komodakis (2018) (CNN) architectures with depth adapted to dataset specifications: a 3-layer CNN (Conv-3) for CIFAR-10/CIFAR-100 synthetic datasets, and a 4-layer CNN (Conv-4) for Tiny-ImageNet and high-resolution ImageNet-1K datasets. More hyperparameter settings are detailed in Appendix B.
@@ 5.2 Comparison with Previous Methods
Tab. 1 demonstrates that the proposed AT-BPTT framework establishes new state-of-the-art performance on three benchmark datasets (CIFAR-10 Krizhevsky et al. (2009), CIFAR-100 Krizhevsky et al. (2009), and Tiny-ImageNet Le and Yang (2015)), outperforming existing inner-loop and outer-loop optimization approaches by a significant margin. Most remarkably, AT-BPTT achieves an average accuracy gain of 6.16% over RaT-BPTT, the SOTA inner-loop DD method. This substantial improvement in low-resolution image distillation stems from the combination of our novel dynamic truncation strategy and adaptive window size mechanism, which collectively enhance the model’s capacity to capture multi-level feature correlations through dynamically adjusted learning trajectories. The visualizations of these synthetic datasets are shown in Appendix F.
@@
Table 1: Comparison with the SOTA baseline dataset distillation methods. Following previous methods, the ConvNet used for distillation are Conv-3 on CIFAR-10 Krizhevsky et al. (2009) and CIFAR-100 Krizhevsky et al. (2009), Conv-4 on Tiny-ImageNet Le and Yang (2015) and ImageNet-1K Russakovsky et al. (2015). Each reported result is the average of 5 experiments. Entries with “-” are absent due to scalability problems.
@@
DSA Zhao et al. (2021) 28.8±0.7 52.1±0.5 60.6±0.5 13.9±0.3 32.3±0.3 42.8±0.4 6.6±0.2 14.4±2.0 1.1±0.7 3.2±0.3
CAFE Wang et al. (2022) 30.3±1.1 46.3±0.6 55.5±0.6 12.9±0.3 27.8±0.3 37.9±0.3 - - - -
MTT Cazenavette et al. (2022) 46.2±0.8 65.4±0.7 71.6±0.2 24.3±0.3 39.7±0.4 47.7±0.2 8.8±0.3 23.2±0.2 - -
TESLA Cui et al. (2023) 48.5±0.8 66.4±0.8 72.6±0.7 24.8±0.4 41.7±0.3 47.9±0.3 - - 7.7±1.0 17.8±0.5
DATM Guo et al. (2024) 46.9±0.5 66.8±0.2 76.1±0.3 27.9±0.2 47.2±0.4 55.0±0.2 17.1±0.3 31.1±0.3 - -
ATT Liu et al. (2024) 48.3±1.0 67.7±0.6 74.5±0.9 26.1±0.5 44.2±0.8 51.2±0.2 11.0±0.4 25.8±0.7 4.7±1.4 8.7±1.0
MCT Zhong et al. (2024) 48.5±0.2 66.0±0.3 72.3±0.3 24.5±0.5 42.5±0.5 46.8±0.2 9.6±0.5 22.6±0.8 - -
NCFM Wang et al. (2025c) 49.5±0.3 71.8±0.3 77.4±0.3 34.4±0.5 48.7±0.3 54.7±0.2 18.2±0.5 26.8±0.6 - -
@@
BPTT Werbos (1990) 49.1±0.6 62.4±0.4 70.5±0.4 21.3±0.6 34.7±0.5 - - - 1.1±0.7 2.3±0.9
FRePO Zhou et al. (2022) 45.6±0.1 63.5±0.1 70.7±0.1 26.3±0.1 41.3±0.1 41.5±0.1 16.9±0.1 22.4±0.1 - -
RCIG Loo et al. (2023) 49.6±1.2 66.8±0.3 - 35.5±0.7 - - 22.4±0.3 - - -
RaT-BPTT Feng et al. (2024) 53.2±0.7 69.4±0.4 75.3±0.3 35.3±0.4 47.5±0.2 50.6±0.2 20.1±0.3 24.4±0.2 5.2±1.1 13.0±0.9
Teddy Yu et al. (2024b) 30.1±1.4 53.0±0.5 66.1±0.4 13.5±0.4 33.4±0.7 49.4±0.5 - - - 34.1±0.8